OpenStudy (anonymous):

I need to find the max and min values by evaluating C=2x+y at critical points with restrictions x>=0, y>=0 and x+y<=8 Is min 0 and max 8 Is max 8 and min 0?

OpenStudy (anonymous):

\[x+y \le8 ...(1)\] consider x+y=8 x=0,y=8,point is B (0,8) y=0,x=8 ,point isA (8,0) put x=0,y=0 in (1) \[0+0 \le 8 ,0\le 8,which~ is~ true. also~ x \ge 0,y \ge 0\] Feasible region is OABO. Draw.  find the value of C at these three points. or plug the values of x and y in C=2x+y and find where it is maximum and where it is minimum.

OpenStudy (anonymous):

so min is 0 and max is 8?